OpenStudy (monica0597):

Identify the correct slope and y intercept of the equation 3x - 3y = 4.

jimthompson5910 (jim_thompson5910):

What do you get when you solve for y?

OpenStudy (monica0597):

-4/3?

OpenStudy (anonymous):

For the equation in the form \(ax + by = c\) Slope is given by: \[Slope = -\frac{x \; \; coefficient}{y \; \; coefficient}\]

jimthompson5910 (jim_thompson5910):

Solve for y to get \[\Large 3x - 3y = 4\] \[\Large -3y = - 3x + 4\] \[\Large y = \frac{- 3x + 4}{-3}\] \[\Large y = \frac{-3x}{-3} + \frac{4}{-3}\] \[\Large y = x - \frac{4}{3}\]

jimthompson5910 (jim_thompson5910):

So what is the slope of \[\Large y = x - \frac{4}{3}\]